# Inertial Coordination Games

### v1

@@ 3   Equilibrium law of motion
Dynamics of threshold beliefs. Recall that agent $i$’s time-$t$ posterior is $N(\mu_{it},\eta_{t}^{2})$. Agent $i$ chooses $a_{it}=1$ at time $t$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, where the belief threshold $\mu_{t}^{*}$ is implicitly defined by the indifference condition
$\mu^{*}_{t}+\mathbb{E}_{\theta\sim N(\mu^{*}_{t},\eta^{2}_{t})}\Big{[}\lambda_{t-1}(\theta)\Big{]}=c$
@@
with boundary $\mu_{1}^{*}:=c-\lambda_{0}$. Note that the dynamics of $(\mu_{t}^{*})_{t}$ are (i) monotone in $t$; (ii) do not depend on the state $\theta$; and (iii) slow down as agents’ beliefs become more precise. Intuitively, when agents are confident about the state, they are confident that their peers have already coordinated on an action, so the indifference condition should then not vary much.
@@
Hence, the limit belief threshold $\mu_{\infty}^{*}$ pins down limit aggregate play at all states $\theta$. In particular, $\lambda_{\infty}(\theta)=RD(\theta)$ a.e. if and only if $\mu_{\infty}^{*}=c-(1/2)$. In words, the risky action is adopted by the whole population in the limit if and only if, in the complete information game in which the fundamental $\theta$ is known, playing the risky action is the best response to the uniform conjecture that $1/2$-proportion of players are playing the risky action. Hence, we call $c-(1/2)$ the risk dominance threshold.
We now turn to two simple but informative examples.
@@
What are we to make of these differing predictions? In what follows we will provide a unified account of when risk-dominance or history-dependence obtains as a function the learning process. It will turn out that Example 1 is a special case of slow learning since posterior precisions grow sub-quadratically (in Example 1, they do not change beyond $t=1$). On the other hand, Example 2 is a special case of fast learning since posterior precisions are super-quadratic (in Example 2, they “jump immediately to infinity” at time $t=1$). [^24] Translated into a static incomplete information game, this correspond to interspersing interim Bayesian updating with rounds of interim deletion of strictly dominated strategies. Appendix D makes this connection precise.
@@ 4   Characterization of limit play
Our main result gives a tight characterization of limit outcomes in terms of the learning rate. All proofs are in Appendix A. We use standard asymptotic notation: $f(t)=O(g(t))$ if there exists constants $\overline{C}<+\infty$ and $T<+\infty$ such that $t\geq T$ implies $f(t)\leq\overline{C}g(t)$. $f(t)=\Omega(g(t))$ if there exists constants $\underline{C}<+\infty$ and $T<+\infty$ such that $t\geq T$ implies $f(t)\geq\underline{C}g(t)$.
@@
    For all $q>2$ and all $\lambda_{0}\in(0,1)$,
    $\eta^{-2}_{t}=\Omega(t^{q})\implies$ $\lambda_{\infty}(\theta)=NRD(\theta)$
@@
Theorem 1 identifies the critical threshold for limit play and, in so doing, develops a tight connection between sub-quadratic growth of posterior precisions and risk-dominance. Economically, it suggests that whether or not coordination succeeds, and how it depends on initial play, depends sharply on the learning rate. We discuss the implications of each part in turn.
Theorem 1 (i) generalizes the global games prediction to the case where learning is sufficiently slow. Thus, our model with an endogenously changing state offers an alternate foundation for equilibrium selection. Moreover, it clarifies that sub-quadratic growth of posterior precision is both necessary and sufficient for risk-dominant selection. Thus, although the ‘vanishing noise limit’ in static global games (Carlsson and Van Damme, 1993; Morris and Shin, 2003) is often interpreted as modelling players with vanishing uncertainty, this actually corresponds to slow learning in our dynamic model with endogenously changing states. [^25] We are grateful to Muhamet Yildiz for pointing out this interpretation.
@@
Parameters: $\theta=0.4$, $\lambda_{0}=0.75$, $c=1$ (not to scale).
@@
Economically, this also implies that shocks to aggregate incentives (which $\lambda_{0}$ can be interpreted as) do not persist when learning is sufficiently slow. Figure 2b (a) illustrates a typical path of aggregate play under slow learning. The true state is $\theta=0.4$ and this governs the location of (heterogeneous) beliefs in the population. The cost of playing the risky action is $c=1$ so under the conjecture that $1/2$-proportion of the population plays the risky action and common-knowledge of the state $\theta=0.4$, the risk-dominant profile is to not play the risky action. Thus, following the initial shock $\lambda_{0}=0.75$, aggregate play converges back to the safe action.
@@
The result in Theorem 1 (ii) that sufficiently fast learning in our model can break risk-dominance is related to the idea that public signals can yield non-risk dominant play even as uncertainty about the fundamentals vanish; see Morris and Shin (2003) Chapter 3.2. In both cases, fast private learning or public signals enable players to quickly approach common knowledge of the fundamental and their peers’ play, thus enabling coordination on the non-risk dominant action. Our main result complements and enhances this insight by providing an explicit dynamic foundation, as well as a precise characterization of how fast learning needs to be for this intuition to hold. This allows us to classify several canonical learning environments and their resulting limit outcomes.
Intuition for Theorem 1. Recall our previous observation that the speed of evolution of the time-$t$ indifferent belief $(\mu_{t}^{*})_{t}$ is inversely related to the speed of learning: when learning is slow enough, $(\mu_{t}^{*})_{t}$ gradually converges to the risk dominant threshold $\mu_{\infty}^{*}=c-(1/2)$ from any initial condition. However, when learning is fast enough, the dynamics of $(\mu_{t}^{*})_{t}$ freeze, and consequently $\mu_{\infty}^{*}$ is bounded away from the risk dominant threshold. The proof of Theorem 1 formalizes this intuition by explicitly constructing fictitious threshold sequences that converge to (non)-risk dominant play, matching them to the appropriate learning rates.
@@
Several corollaries follow relatively straightforwardly from Theorem 1.
@@
This gives an explicitly dynamic foundation for the gap between almost-common knowledge and common knowledge of payoffs in static incomplete information games. Indeed, Corollary 1 suggests that the Gaussian form of private signals is not necessary since, for (i) an arbitrary noise structure (with bounded variance); and (ii) “sufficiently smooth priors”, as $T$ grows large the distribution over posterior beliefs after $T$ grows close to the normal distribution by the Bernstein-von Mises Theorem (Van der Vaart, 2000).
@@
    (Social learning) At each time $t\geq 2$, player $i$ exchanges information with player $j$. Thus, defining $\Sigma(\sigma)$ to be: $\sigma_{1}^{2}:=\sigma^{2}$ and for all $t\geq 2$, $\sigma_{t}^{2}:=\eta_{t-1}^{2}$, we have that for any $\lambda_{0}$, there exists some $\underline{\sigma}>0$ such that under $\Sigma(\sigma)$ for any $\sigma<\underline{\sigma}$, there exists an open interval of states at which limit aggregate play is the non-risk dominant action.
@@
    Extreme initial play induces extreme limit play. For any $|\lambda_{0}-(1/2)|\leq|\lambda_{0}^{\prime}-(1/2)|$ and any $\Sigma$,
@@
    Monotonicity in learning speeds. For any $\Sigma\geq\Sigma^{\prime}$,
@@
    Then, for any $s\geq s^{\prime}$ and $\sigma_{s}^{2}\geq\sigma_{s^{\prime}}^{2}$,
@@
In particular, for any threshold $\mu^{*}$ between $c-\lambda_{0}$ and $c-(1/2)$, there exists some learning rate $\Sigma\geq 0$ such that:
@@
The intuition behind Proposition 2 is that for any time $t$, we can arbitrarily restrict the evolution of the system by choosing a sufficiently precise signal for time $t$. This verifies that any level of coordination on the non-risk dominant action can be rationalized by some sequence of signal noises. [^27] Recall that $\mu_{1}^{*}=c-\lambda_{0}$ and $\mu_{t}^{*}$ evolves monotonically away. Together with our comparative static on learning speeds (Proposition 1 (ii)) this also implies $\lim_{\Sigma\downarrow 0}\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\mathbbm{1}(\theta\geq c-\lambda_{0})$—that is, in the vanishing noise limit along the whole path, limit play converges to that of the complete environment with all players best-responding to $\lambda_{0}$.
@@ 5   Transition dynamics
(a) $\sigma=1$
(b) $\sigma=0.01$
@@ 6   Concluding Remarks
We have developed a simple but novel model of dynamic coordination with (i) Bayesian learning about fundamental uncertainty; and (ii) adaptive best-response dynamics. This has allowed us to combine and unify insights from the literature on best-response dynamics in complete-information coordination games (Crawford, 1995) and global games (Carlsson and Van Damme, 1993; Morris and Shin, 2003). Our model serves as a natural description of many economic environments with both shrinking uncertainty about fundamentals (e.g., solvency of a bank, strength of the economy, quality of a networked product) as well as an endogenous time-varying state which is influenced by aggregate past play.
Within our environment, we uncovered a tight connection between learning rates and limit play (Theorem 1) and established sub/super-quadratic reduction in posterior variance as the critical threshold at which limit play undergoes a shift away from risk-dominance. We also show that, subject to initial conditions, any shift from risk-dominance can be attained by some learning rate. Moreover, in the sub-quadratic regime, we further identified qualitatively different behavior in the path of aggregate play (Proposition 3): when learning is slow, transition dynamics are gradual; when learning is fast, sudden.
@@ Appendix A Omitted proofs
Section A.1 contains several useful lemmas that characterize the dynamics of our model. Sections A.2-A.5 collect all remaining proofs. We sometimes assume $\lambda_{0}>\frac{1}{2}$ without loss of generality because symmetric arguments apply to $\lambda_{0}<\frac{1}{2}$.
@@ A.1  System Dynamics
At each time $t\geq 1$ and for each player $i$, $a_{it}=1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, where $\mu_{1}^{*}:=c-\lambda_{0}$ and for all $t>1$, $\mu_{t}^{*}$ is the unique solution to the nonlinear difference equation
@@
Assume $\mu_{t-1}^{*}=c-(1/2)$. Since $\mu_{t}^{*}:=c-(1/2)$ solves ($\mu^{*}$-LoM) and it has a unique solution, $c-(1/2)$ is a steady state. To see that it’s the unique steady state, assume $\mu_{t-1}^{*}=\mu_{t}^{*}$. Then, $\mu_{t}^{*}=c-(1/2)$, as desired.
Next, fix any $t\in\mathcal{T}$. By ($\mu^{*}$-LoM), we have
@@
Note that $\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}$ is non-increasing in $t$. Wlog, assume $\lambda_{0}>\frac{1}{2}$. Then, $(c-\mu_{t+1}^{*})$ decreasing in $t$ implies $\Phi^{-1}(c-\mu_{t+1}^{*})>0$ is decreasing in $t$, and therefore $|\Phi^{-1}(c-\mu_{t+1}^{*})|$ is decreasing in $t$. Hence, $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is also decreasing in $t$.
@@
$0<\frac{\theta-\mu_{\infty}^{*}}{\eta_{t}}<\frac{\theta-\mu_{t}^{*}}{\eta_{t}}.$
@@
$\gamma_{t-1}\geq\gamma_{t}\iff\Phi(\gamma_{t-1})\geq\Phi(\gamma_{t})\iff\gamma_{t}\geq 0$
Wlog, assume If $\lambda_{0}>(1/2)$, then $\gamma_{1}>0$. Suppose for a contradiction that $\gamma_{2}<0$: then $\gamma_{1}<\gamma_{2}<0$, which is a contradiction since $\gamma_{1}>0$ by definition. Hence, $\gamma_{1}\geq\gamma_{2}\geq 0$. By induction, $(\gamma_{t})_{t}$ is a monotone decreasing sequence bounded below by 00, so it converges: $\gamma_{t}\downarrow\gamma_{\infty}\geq 0$, as desired. ∎
@@
By an exactly analogous argument as in Lemma 4(ii), note that $x>0$ implies $x\geq F_{A}(x)\geq 0$, which implies $\Phi(x)-\Phi(F_{A}(x))\in[0,\Phi(x)-(1/2)]$ and hence
@@
Note that ($\gamma$-LoM) states: $\gamma_{t}=F_{A_{t}}(\gamma_{t-1})$.
@@
Note that ($\gamma$-LoM) states: $F_{A_{\gamma^{*}}(\gamma_{t-1})}(\gamma_{t-1})=(\gamma^{*}+\gamma_{t-1})/2$.
@@
Then, $A_{\gamma_{1}}(\gamma_{1})=0$ and $A_{\gamma_{1}}$ is decreasing on $[0,\gamma_{1}]$. Thus, for any fixed $\gamma^{*}\in(0,\gamma_{1})$,
$A_{\gamma_{1}}(x)\uparrow A_{\gamma_{1}}(\gamma^{*})=:\bar{A}^{*}$
@@
Note that
@@
Note that ($\gamma$-LoM) states: $F_{A_{\gamma_{1}}(x)}(\gamma_{1})=x$. With these lemmas in hand, we can understand how $(\gamma_{t})_{t}$ changes as a function of $(A_{t})_{t}$.

### v2

@@ 3   Equilibrium law of motion
Dynamics of threshold beliefs. Recall that agent $i$’s time-$t$ posterior is $N(\mu_{it},\eta_{t}^{2})$. Agent $i$ chooses the risky action $a_{it}=1$ at time $t$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, where the belief threshold $\mu_{t}^{*}$ is implicitly defined by the indifference condition
$\mu^{*}_{t}+\mathbb{E}_{\theta\sim N(\mu^{*}_{t},\eta^{2}_{t})}\Big{[}\lambda_{t-1}(\theta)\Big{]}=c.$
@@
with boundary $\mu_{1}^{*}:=c-\lambda_{0}$. Note that the dynamics of $(\mu_{t}^{*})_{t}$ are (i) monotone in $t$; (ii) do not depend on the state $\theta$; and (iii) slow down as agents’ beliefs become more precise. Intuitively, when agents are confident about the state, they are confident that their peers have already coordinated on an action, so the indifference condition does not then not vary much.
@@
Hence, the limit belief threshold $\mu_{\infty}^{*}$ pins down limit aggregate play at state $\theta$. In particular, $\lambda_{\infty}(\theta)=RD(\theta)$ a.e. if and only if $\mu_{\infty}^{*}=c-(1/2)$. In words, the risky action is adopted by the whole population in the limit if and only if, in the complete information game in which the fundamental $\theta$ is known, playing the risky action is the best response to the uniform conjecture that $1/2$-proportion of players are playing the risky action. Hence, we call $c-(1/2)$ the risk dominance threshold.
We now walk through two simple but illustrative examples.
@@
What are we to make of these differing predictions? In what follows we offer a unified account of when risk-dominance or history-dependence obtains as a function of the learning process. It will turn out that Example 1 is a special case of slow learning since posterior precisions grow sub-quadratically (in Example 1, they do not change beyond $t=1$). On the other hand, Example 2 is a special case of fast learning since posterior precisions are super-quadratic (in Example 2, they “jump immediately to infinity” at time $t=1$). [^27] Translated into a static incomplete information game, this correspond to interspersing interim Bayesian updating with rounds of interim deletion of strictly dominated strategies. Online Appendix V makes this connection precise.
@@ 4   Characterization of limit play
Our main result gives a tight characterization of limit outcomes in terms of the learning rate. All proofs are in Appendix A. To state our main result, we will use standard asymptotic notation: $f(t)=O(g(t))$ if there exist constants $\overline{C}<+\infty$ and $T<+\infty$ such that $t\geq T$ implies $f(t)\leq\overline{C}g(t)$. $f(t)=\Omega(g(t))$ if there exist constants $\underline{C}<+\infty$ and $T<+\infty$ such that $t\geq T$ implies $f(t)\geq\underline{C}g(t)$.
@@
    For all $\epsilon>0$ and all $\lambda_{0}\in(0,1)$,
    $\eta^{-2}_{t}=\Omega(t^{2+\epsilon})\implies$ $\lambda_{\infty}(\theta)=NRD(\theta)$
@@
Theorem 1 identifies the critical threshold for limit play and, in so doing, develops a tight connection between sub-quadratic growth of posterior precisions and risk-dominance. Economically, it suggests that in the presence of (a little) inertia, whether and how shocks propagate to shape coordination outcomes depends sharply on learning speeds. We discuss the implications of each part in turn.
Theorem 1 (i) generalizes the global games prediction to the case where learning is sufficiently slow. Thus, our model with an endogenously changing state offers an alternate foundation for equilibrium selection. Moreover, it clarifies that sub-quadratic growth of posterior precision under private learning is both necessary and sufficient for risk-dominant selection. Thus, although the ‘vanishing noise limit’ in static global games (Carlsson and Van Damme, 1993; Morris and Shin, 2003) is often interpreted as modeling players with vanishing uncertainty, this actually corresponds to slow learning in our dynamic model with endogenously changing states. [^28] We are grateful to Muhamet Yildiz for pointing out this interpretation.
@@
Parameters: $\theta=0.4$, $\lambda_{0}=0.75$, $c=1$
@@
Economically, this also implies that shocks to aggregate incentives (which we can interpret $\lambda_{0}$ as) do not persist when learning is sufficiently slow. Figure 2b (a) illustrates a possible path of aggregate play under slow learning. [^29] Notice that $(\lambda_{t})_{t}$ ‘jumps’ toward the risk-dominant profile. In Section 5 we identify conditions under which the transition to the risk-dominant action is ‘gradual’ versus ‘sudden’. The true state is $\theta=0.4$ and this governs the location of (heterogeneous) beliefs in the population. The cost of playing the risky action is $c=1$ so under the conjecture that $1/2$-proportion of the population plays the risky action and common-knowledge of the state $\theta=0.4$, the risk-dominant profile is to not play the risky action. Thus, following the initial shock $\lambda_{0}=0.75$, aggregate play converges back to the safe action.
@@
Putting parts (i) and (ii) of Theorem 1 together offers simple but novel economic implications for how the ex-ante joint distribution over shocks $(\lambda_{0})$ and fundamentals ($\theta$) interact with learning speeds to shape limit outcomes. Figure 3b depicts the regions of realizations of shocks and fundamentals under which shocks propagate so the limit play is the risky action (red region), or fizzle out so limit play is the safe action (blue region). Panel (a) depicts this for the case of sub-quadratic learning. As Theorem 1 (i) establishes, the risk-dominant boundary is selected so limit play is history independent, and does not depend on the shock’s realization. Panel (b) depicts the case of super-quadratic learning which rotates the boundary clockwise around the point $(1/2,c-1/2)$. [^30] We have depicted this so that the state selects initial-play dominance i.e., $\lambda_{\infty}=1\iff\theta\geq c-\lambda_{0}$. We will later establish (Proposition 2 that initial-play dominance and risk-dominance span possible limit outcomes implementable via any learning speed $(\sigma_{t})_{t}$. More generally, for a given super-quadratic learning rate, this will lead to a (potentially non-linear) boundary between these two extremes. Hence, when learning is fast, the regime might now survive when the state is weak as long as the initial shock is small (region $A$); on the other hand, the regime might now also fail when the state is strong when the initial shock is large (region $B$).
Figure 3: How learning speeds shape limit play
(a) Sub-quadratic
(b) Super-quadratic
That sufficiently fast learning can break risk-dominance is related to the idea that public signals can yield non-risk dominant play even as uncertainty about the fundamentals vanish; see Morris and Shin (2003) Chapter 3.2. In both cases, fast private learning or public signals enable players to quickly approach common knowledge of the fundamental and their peers’ play, thus enabling coordination on the non-risk dominant action. Our result complements and enhances this insight by providing an explicit dynamic foundation, as well as a precise characterization of how fast learning needs to be for this intuition to hold. This allows us to classify several canonical learning environments and their resulting limit outcomes.
Intuition for Theorem 1. Recall our previous observation that the speed of evolution of the time-$t$ indifferent belief $(\mu_{t}^{*})_{t}$ is inversely related to the speed of learning: when learning is slow enough, $(\mu_{t}^{*})_{t}$ gradually converges to the risk dominant threshold $\mu_{\infty}^{*}=c-(1/2)$ from any initial condition. However, when learning is fast enough, the dynamics of $(\mu_{t}^{*})_{t}$ freeze, and consequently $\mu_{\infty}^{*}$ is bounded away from the risk dominant threshold. The proof of Theorem 1 formalizes this intuition by explicitly constructing fictitious threshold sequences that converge to (non)-risk dominant play and matching them to the appropriate learning rates.
@@
Several corollaries follow from Theorem 1.
@@
This gives a dynamic perspective on the gap between almost-common knowledge and common knowledge of payoffs in static incomplete information games.
@@
    (Social learning) At each time $t\geq 2$, player $i$ exchanges information with a randomly matched player $j$. That is, in the first period, each play receives a private signal $\sigma^{2}_{1}=\sigma^{2}$ and in all subsequent periods, players’ posterior precisions double: $\sigma^{2}_{t}=\eta^{2}_{t-1}$. Then the non-risk dominant action is played on a positive measure of states. In fact, we will later show that if $\sigma$ is small, limit play converges to initial play dominance where $\lambda_{\infty}(\theta)=1$ if and only if $\theta\geq\lambda_{0}$.
@@
    Extreme initial play induces extreme limit play. For any $|\lambda_{0}-(1/2)|\leq|\lambda_{0}^{\prime}-(1/2)|$ and any $\Sigma$, the limit threshold is further from the risk-dominance threshold when initial play is more extreme:
@@
    Monotonicity in learning speeds. For any $\Sigma\geq\Sigma^{\prime}$, the limit threshold is further from the risk-dominance threshold when learning is quicker:
@@
    Then, the limit threshold is further from the risk-dominance threshold when informative signals are front-loaded: for any $s\geq s^{\prime}$ and $\sigma_{s}^{2}\geq\sigma_{s^{\prime}}^{2}$,
@@
In particular, for any threshold $\mu^{*}$ between $c-\lambda_{0}$ and $c-(1/2)$, there exists some learning process $\Sigma\geq 0$ such that:
@@
The intuition behind Proposition 2 is that for any time $t$, we can arbitrarily restrict the evolution of the system by choosing a sufficiently precise signal for time $t$. That is, any level of coordination on the non-risk dominant action can be rationalized by some sequence of signal noises. [^32] Recall that $\mu_{1}^{*}=c-\lambda_{0}$ and $\mu_{t}^{*}$ evolves monotonically away. Combined with our comparative static on learning speeds, this also implies $\lim_{\Sigma\downarrow 0}\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\mathbbm{1}(\theta\geq c-\lambda_{0})$—that is, in the vanishing noise limit along the whole path, limit play converges to that of the complete environment with all players best-responding to $\lambda_{0}$.
@@ 5   Transition dynamics
(a) $\sigma=1$: gradual transition
(b) $\sigma=0.01$: sudden transition
@@ 6   Concluding Remarks
We have developed a simple but novel model of inertial coordination games. This allowed us to combine and unify insights from the literature on best-response dynamics in complete-information coordination games (Crawford, 1995) and global games (Carlsson and Van Damme, 1993). We think that inertial coordination games serve as a natural description of many ‘real world’ coordination environments with both shrinking uncertainty about fundamentals (e.g., solvency of a bank, strength of a currency regime, quality of a networked product) as well as an endogenously changing state shaped by past play (e.g., past withdrawals, short positions, past sales). At the same time, we also showed that inertial coordination games share much of the game-theoretic logic underpinning coordination and learning in games. [^34] We show in Appendix B that for a fixed speed of learning (as a function of physical time), limit equilibrium predictions coincide as inertia vanishes (gap between time periods shrink). In this regard, inertial coordination games generalizes the setting with contemporaneous coordination motives. Nonetheless for any fixed level of inertia, the sub/super-quadratic boundary continues to hold. In this regard, we offer a different perspective on ‘history versus expectations’ (Krugman, 1991) in economic modeling—in the presence of inertia, expectations about history shape coordination outcomes. In particular, the speed at which players grow confident is a crucial determinant for whether coordination succeeds—the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically (Theorem 1). This delivered a suite of rich predictions of whether and to what extent initial shocks can propagate. In the super-quadratic regime, we showed that any profile of limit play between risk- and initial play-dominance can be rationalized by some learning speed. In the sub-quadratic regime, we further identified qualitatively different behavior in the path of aggregate play (Proposition 3): when learning is slow, transition dynamics are gradual; when learning is fast, sudden.
@@ Appendix A Omitted proofs
Section A.1 contains several useful lemmas that characterize the dynamics of our model. Sections A.2-A.5 collect all remaining proofs. We sometimes assume $\lambda_{0}>\frac{1}{2}$ without loss of generality because symmetric arguments apply to $\lambda_{0}<\frac{1}{2}$. In particular, the proofs of Theorem 1 and Proposition 2 consist of several lemmas, as described in Figure 8.
Figure 8: Roadmap for proofs of Theorem 1 and Proposition 2
@@ A.1  System Dynamics
At each time $t\geq 1$ and for each player $i\in[0,1]$, $a_{it}=1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, where $\mu_{1}^{*}:=c-\lambda_{0}$ and for all $t>1$, $\mu_{t}^{*}$ is the unique solution to the nonlinear difference equation
@@
Assume $\mu_{t-1}^{*}=c-(1/2)$. Since $\mu_{t}^{*}:=c-(1/2)$ solves ($\mu^{*}$-LoM) and it has a unique solution, $c-(1/2)$ is a steady state. To see that it’s the unique steady state, assume $\mu_{t-1}^{*}=\mu_{t}^{*}$. Then, ($\mu^{*}$-LoM) reduces to $\mu_{t}^{*}+\Phi(0)=c$ and hence $\mu_{t}^{*}=c-(1/2)$, as desired.
Next, fix any $t\in\mathcal{T}$. We have
@@
Note that $\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}$ is non-increasing in $t$. First, assume $\lambda_{0}>\frac{1}{2}$. Then, $(c-\mu_{t+1}^{*})>\frac{1}{2}$ decreasing in $t$ implies $\Phi^{-1}(c-\mu_{t+1}^{*})>0$ is decreasing in $t$, and therefore $|\Phi^{-1}(c-\mu_{t+1}^{*})|$ is decreasing in $t$. Hence, $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is also decreasing in $t$.
Next, assume $\lambda_{0}<\frac{1}{2}$. Then, $(c-\mu_{t+1}^{*})<\frac{1}{2}$ increasing in $t$ implies $\Phi^{-1}(c-\mu_{t+1}^{*})<0$ increasing in $t$, and therefore $|\Phi^{-1}(c-\mu_{t+1}^{*})|$ is decreasing in $t$. Hence, $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is also decreasing in $t$.
@@
$0<\frac{\theta-\mu_{\infty}^{*}}{\eta_{t}}<\frac{\theta-\mu_{t}^{*}}{\eta_{t}}.$ ($\mu_{t}^{*}<\mu_{\infty}^{*}$)
@@
$\gamma_{t-1}\geq\gamma_{t}\iff\Phi(\gamma_{t-1})\geq\Phi(\gamma_{t})\iff\gamma_{t}\geq 0$ (From ($\gamma$-LoM))
Wlog, assume $\lambda_{0}>(1/2)$, then $\gamma_{1}>0$. Suppose for a contradiction that $\gamma_{2}<0$: then $\gamma_{1}<\gamma_{2}<0$, which is a contradiction since $\gamma_{1}>0$ by definition. Hence, $\gamma_{1}\geq\gamma_{2}\geq 0$. By induction, $(\gamma_{t})_{t}$ is a monotone decreasing sequence bounded below by 00, so it converges: $\gamma_{t}\downarrow\gamma_{\infty}\geq 0$, as desired. ∎
Next, we study the function $F_{A}$ such that ($\gamma$-LoM) implies: $\gamma_{t}=F_{A_{t}}(\gamma_{t-1})$.
@@
Next, we follow the argument of Lemma 4(ii). Assume $x>0$, and suppose for a contradiction that $x<F_{A}(x)$. Then, $AF_{A}(x)=\Phi(x)-\Phi(F_{A}(x))<0$ and $A>0$ implies $F_{A}(x)<0<x$, a contradiction. Hence, $x\geq F_{A}(x)\geq 0$, which implies $\Phi(x)-\Phi(F_{A}(x))\in[0,\Phi(x)-(1/2)]$ and hence
@@
Next, we study the function $A_{\gamma^{*}}$ such that ($\gamma$-LoM) implies: $F_{A_{\gamma^{*}}(\gamma_{t-1})}(\gamma_{t-1})=(\gamma^{*}+\gamma_{t-1})/2$.
@@
Finally, we study the function $A_{\gamma_{1}}$ such that ($\gamma$-LoM) implies: $F_{A_{\gamma_{1}}(x)}(\gamma_{1})=x$.
@@
Then, $A_{\gamma_{1}}(\gamma_{1})=0$ and $A_{\gamma_{1}}(x)$ is decreasing on $x\in[0,\gamma_{1}]$. Thus, for any fixed $\gamma^{*}\in(0,\gamma_{1})$,
$A_{\gamma_{1}}(x)\uparrow A_{\gamma_{1}}(\gamma^{*})=:\bar{A}^{*}>0$
@@
$A_{\gamma_{1}}(\gamma_{1})=0$ follows immediately from the definition. Note that $A_{\gamma_{1}}(x)$ is continuous and differentiable on $x\in[0,\gamma_{1}]$.
@@
With these lemmas in hand, we can understand how $(\gamma_{t})_{t}$ changes as a function of $(A_{t})_{t}$.